OpenStudy (anonymous):

L(T)= (1/2)^t Explain why this is a decreasing function. Can anyone explain this?

OpenStudy (anonymous):

Meant to be L(t)

OpenStudy (binary3i):

because 2^t will keep on increasing. you tell me which noumber is greater 1/64 or 1/2

OpenStudy (kinggeorge):

If you're familiar with derivatives, that's the best way to show it.\[\large L'(t)=\ln\left(\frac{1}{2}\right)\cdot\left(\frac{1}{2}\right)^t\]Since \(\ln(1/2)<0\) this derivative is negative. This implies that the function is always decreasing.

OpenStudy (anonymous):

Thank you :)

OpenStudy (kinggeorge):

You're welcome.